# On the Fundamental Limit of Stochastic Gradient Identification Algorithm Under Non-Persistent Excitation

### v1

@@ 2 Problem Formulation
$\varphi_{n}^{\top}=[y_{n-1}^{\top},\cdots,y_{n-p+1}^{\top},$ (9)
$u_{n-1}^{\top},\cdots,$ $u_{n-q+1}^{\top},$
$\hat{w}_{n-1}^{\top},\cdots,\hat{w}_{n-r+1}^{\top}],$
@@
Through a counterexample inspired by lai1982, Chen and Guo show that if the condition number grows strictly faster than $O(\log r_{n})$, the SG algorithm can be made to fail.
@@ 3 Mathematical Framework
In this section, to ensure our theoretical framework is sufficiently general to address a broad class of problems, we let $0\leq A_{n}\leq I$ denote general positive semi-definite symmetric matrices, which admit a decomposition of the form
@@
Now, introduce a nonnegative real sequence $\mu_{n}\geq 0$ called the weights and define a weighted sum $S_{ik}$ of the matrices $A_{n}$ over the interval $[k,i]$:
@@
We estimate the quadratic form:
@@
Thus, we obtain the following inequality relating $\|x_{k}\|$ and $\|x_{i}\|$:
@@
$\leq\left[\sqrt{\max_{k\leq j<i}\mu_{j}}+\sqrt{\sum_{j=k}^{i-1}\mu_{j}B_{jk}}\right](\|x_{k}\|^{2}-\|x_{i}\|^{2}).$
@@ 4 Convergence Analysis
where $t_{k}$ is defined similarly, $S_{t_{k}t_{k-1}}=\sum_{i=t_{k-1}}^{t_{k}-1}\varphi_{i}\varphi_{i}^{\top}$, and $D_{k}=r_{t_{k}-1}\left(\log r_{t_{k}-1}-\log r_{t_{k-1}-1}\right)-(r_{t_{k}-1}-r_{t_{k-1}-1})$.

### v2

@@ 2 Problem Formulation
$\varphi_{n}^{\top}=[y_{n}^{\top},\cdots,y_{n-p+1}^{\top},$ (9)
$u_{n}^{\top},\cdots,$ $u_{n-q+1}^{\top},$
$\hat{w}_{n}^{\top},\cdots,\hat{w}_{n-r+1}^{\top}],$
@@
Remark 1
A linear regression model can be viewed as an ARMAX model with $p=r=0,q=1$.
@@
Through a counterexample, Chen and Guo show that if the condition number grows strictly faster than $O(\log r_{n})$, the SG algorithm can be made to fail.
@@ 3 Mathematical Framework
In this section, we let $0\leq A_{n}\leq I$ denote general symmetric matrices of rank 1, which admit a decomposition of the form
@@
Now, introduce a nonnegative real sequence $\mu_{n}\geq 0$ called the weights and define a weighted sum $S_{ik}$ of the matrices $A_{n}$ over the interval $[k,i)$:
@@
We compute the quadratic form:
@@
Combining (34), (36), and (37), we obtain the following inequality relating $\|x_{k}\|$ and $\|x_{i}\|$:
@@
$\leq\left[\sqrt{\max_{k\leq j<i}\mu_{j}}+\sqrt{\sum_{j=k}^{i-1}\mu_{j}B_{jk}}\right]^{2}(\|x_{k}\|^{2}-\|x_{i}\|^{2}).$
@@ 4 Convergence Analysis
where $t_{k}$ is defined similarly, $S_{t_{k}t_{k-1}}=\sum_{i=t_{k-1}}^{t_{k}-1}\varphi_{i}\varphi_{i}^{\top}$, and $D_{k}=r_{t_{k}-1}\left(\log r_{t_{k}-1}-\log r_{t_{k-1}-1}\right)+r_{t_{k-1}-1}$.